OpenStudy (anonymous):

In a right triangle, the length of the hypotenuse is point five centimeters. If the length of one leg is zero point three centimeters, what is the length of the other leg? point two centimeters point four centimeters point six centimeters point eight centimeters

OpenStudy (anonymous):

@HelpBlahBlahBlah

OpenStudy (anonymous):

@mathlover2014

OpenStudy (anonymous):

I think it is d, but idk.... can you explain why?

OpenStudy (anonymous):

HELP!

OpenStudy (anonymous):

@mtbender74

OpenStudy (anonymous):

the pythagorean theorem is \[a^{2} + b^{2} = c^{2}\] where a and b are the legs, and c is the hypotenuse you have a=0.3, b=?, c=0.5 plug in those values and solve for b

OpenStudy (anonymous):

So 0.2?

OpenStudy (anonymous):

not quite... \[(0.3)^{2} + b^{2} = (0.5)^{2}\] \[0.09 + b^{2} = 0.25\] \[b^{2} = 0.25 - 0.09\] so what does b =?

OpenStudy (anonymous):

0.4

OpenStudy (anonymous):

much better :)